# Large Language Models for Orchestrating Bimanual Robots

### v1

@@ IV-C Results
In this section, we report the results of the LABOR agent in controlling the NICOL robot to solve the above tasks, including success rate and failure analysis. To evaluate the effectiveness, we repeat each task 100 times and for each time a variant of the task is randomly chosen. The success rates and failure rates in three aspects of LABOR when using GPT-4 as the LLM on these tasks are listed in Table II.
TABLE II: Success and Failure rates of NICOL on three everyday tasks
Task ControlScissors ServeWater HoldBowl
Success 79% 89% 86%
Failure Spatial 12% 5% 4%
Temporal 1% 6% 1%
Skill 8% 0% 9%
@@
It is observed that the LABOR agent with GPT-4 demonstrates extraordinary performance in orchestrating bimanual control in long-horizon manipulation tasks. Regarding the properties of bimanual manipulation and the LLM itself, we found that the failure cases mainly come from three aspects including Temporal Coordination, Spatial Coordination, and Skill Association. In this sense, Temporal failure refers to the fact that the LLM fails to generate skills in the correct order, Spatial failure refers that the LLM mishandles the spatial relationships in the operation, such as the relationship between two objects or between the robot’s hand and an object, and Skill failure refers to the understanding deviance of the LLM in the correlation between the designed skills and the task requirements. To simplify the analysis process, failures are counted based on the error type that occurred the first time.
It was observed that GPT-4 encountered some difficulties with spatial reasoning (i.e., moving in a certain direction to open or close the scissors) in the ControlScissors task. Whereas, the main errors in the ServeWater task came from temporal and spatial coordination, such as executing pouring before moving the blue cup over the yellow cup, executing grasping when the selected hand had already grasped another object, executing grasping when the object that was outside of the selected hand’s workspace, etc. Moreover, some skill comprehension biases happened in the ControlScissors and HoldBowl tasks, such as moving the scissors handle with the hand currently pressing on it, executing grasping the bowl with a single hand, and so on.
In contrast, when implementing LABOR with GPT-3.5, it only achieved a 49% success rate in the ControlScisssor task, which might be due to the task requiring fewer planning steps. Apart from that, it demonstrates its inability to accomplish either ServeWater or HoldBowl tasks. The failures observed in these tasks couldn’t be neatly categorized into three aspects, as in such long-horizon tasks, the GPT-3.5 model’s understanding of the skills and its temporal and spatial coordination was incorrect at the first step. In addition, it is not sensitive enough to the information feedback by the environment, causing it to stay stuck in repeating wrong commands.
In addition to the success rate, we also counted the lengths of the skill chains generated by GPT-4 in all cases where the task was successfully solved, and compared them with the optimal human-designed skill chain lengths, as shown in Fig. 7. Since there are variants in the ServeWater and HoldBowl quests that single-hand sequential operation, the statistics in these two tasks are reported separately based on the position in which the two objects are located, i.e., Same Side and Two Sides. The number of timesteps indicated by the dashed lines is considered to be the optimal lower bound for the number of timesteps required. We can observe that there are many cases in ServeWater and ControlScissors tasks, where an optimal skill chain can be achieved that is comparable to the human level, while it is pretty close in the HoldBowl task. To ensure the safety of the two-handed operation, the GPT-4 generates some waits for single hands thus bringing about an increase in the length of the chain. At the same time, humans understand the task well enough, thus allowing for the most efficient utilization of both hands. Notably, due to GPT-4’s strong in-context reasoning and understanding ability, sometimes long skill chains can still ensure task accomplishment in the ServeWater and HoldBowl tasks.
In summary, the LABOR agent with GPT-4 shows remarkable performance in coordinating the NICOL robot in long-horizon tasks, not only in terms of the success rate but also in outputting skill chains comparable to the human level.
@@ V Conclusions
In this work, we introduce LABOR, an agent that utilizes an LLM to effectively orchestrate bimanual control when performing robotic manipulation tasks. Experimental results on the NICOL robot for three typical everyday tasks showcase the effectiveness of the LABOR agent.
Limitations and Future Work. Similar to other LLM-based methods, the LABOR agent is dependent on the performance of LLM like the GPT-4 model, requiring its reasoning and understanding of the correlations of primitive skills and task solutions. Besides, since the agent is currently solely running in the simulated environment instead of the real world, some potential differences in environmental dynamics might not be sufficiently analyzed or considered, which brings some unknowns in transferring the agent. In future work, the learning-based skills combined with vision-based models will be developed to reinforce the abilities of LABOR agent in bimanual manipulation. As skills expand, the LABOR agent should be able to control bimanual robots to perform more complex tasks. Furthermore, it is a promising direction to include obstacle detection and collision-free motion planning models in dealing with rich-contact tasks.
@@ VI ACKNOWLEDGMENT
The authors would like to thank Jan-Gerrit Habekost and Lennart Clasmeier for their work and support on the NICOL robot in the simulation.

### v2

@@ IV-C Results
In this section, we report the results of the LABOR agent in controlling the NICOL robot to solve the above tasks, including success rate and failure analysis. To evaluate the effectiveness, we repeated 40 tasks for each class, and for each time a variant of the task was randomly chosen. The success rates and failure rates in three aspects of LABOR when using GPT-4o as the LLM on these tasks are listed in Table I. It can be observed that, compared with the baseline method, the LABOR agent demonstrates higher performance in orchestrating bimanual control in long-horizon manipulation tasks, especially in the ServeFruit task.
@@
Regarding the properties of bimanual manipulation and the LLM itself, we found that the failure cases mainly come from two aspects namely Temporal Coordination and Spatial Coordination. In this sense, Temporal failure refers to the fact that the LLM fails to generate skills in the correct order or neglects some processes during the task execution, while Spatial failure refers that the LLM mishandles the spatial relationships in the operation, such as the relationship between two objects or between the robot’s hand and an object.
By checking the plans generated by LLMs in each failed task, we found that in the ServeWater task, the Temporal failures mainly occurred when the pouring water process was not executed during the task accomplishment, while Spatial failures mainly on that such process happened without moving above the yellow cup, causing the water not being transferred to another cup. Similarly in the ServeFruit task, Spatial failures mainly come from the incorrect releasing process of the apple or banana, while the Temporal failures are mainly due to the lack of such correct processes, causing the bowl to be served to the target point without fruits or only one single fruit.
TABLE I: Experiment results on two classes of bimanual robotic tasks showcase that our method outperforms baseline in terms of success rate.
In addition to this, we also see a few cases in Baseline agent where the LLM tries to perform a transfer of an object through two hands through the overlap area, instead of performing a manipulation about two objects through the overlap area. For example, in the Servewater task, the LLM tries to move the blue cup to another area to pour the water and then transfer it back through the overlap area, which is an interesting phenomenon, but unfortunately due to the long number of steps required for this kind of reasoning, the LLM interrupts prematurely resulting in the task not proceeding to the last step. In contrast, for the LABOR agent, the LLM innovatively uses the skill of push_to to push the blue cup to the overlap area, which is feasible, but then the LLM ignores the fact that the cup is not held by the hand (requiring move_and_grasp process), failing in the task eventually.
Excluding these failure cases, one can easily observe that the LABOR agent exceeds the success rate of the Baseline agent in both long-horizon bimanual tasks, especially the ServeFruit task, demonstrating its outstanding efficacy in coordinating two hands in complex tasks.
IV-D Discussion
Although the guiding prompt is structured to coordinate two-hand control while allowing each hand to operate independently, it inadvertently biases the LLM towards using both hands as much as possible. Due to the generic skills we designed for the NICOL robot and the flexibility provided by the overlap area, the LLM agents occasionally exhibit unexpected behaviors. For instance, when two cups are in different areas, the most efficient approach would be to lift both cups to the overlap area to pour out the water and prepare for future actions. However, the LLM agent might instead transfer the object from the left hand to the right across the overlap area, pouring out the water in a different area.
Experimental results show that these behaviors still achieve the desired outcomes. For example, when fruits are located in different areas, the most efficient method is to lift and release the fruit on the same side as the bowl and push the bowl to the overlap area to add another fruit. However, the LLM occasionally uses the overlap area as a transition point to transfer the fruit, which, although accomplishing the task, slows down progress and incurs a higher likelihood of task failure. While we anticipated some anomalies when designing the skills, these unanticipated behaviors provide valuable insights for developing more complex skills in future real-world applications. For example, designing fine-grained precondition detection mechanisms for each skill or describing these conditions accurately in prompts so that LLMs can better understand the skill.
@@ V Conclusions
In this work, we introduce LABOR, an agent that utilizes an LLM to effectively orchestrate bimanual control when performing humanoid manipulation tasks. Under the spatio-temporal control patterns designed in guiding prompts, the LABOR agent decomposes the task into several stages involving uncoordinated and coordinated steps, according to the spatial relationships and task requirement. Based on the generated plan, the LLM then iteratively generates action steps based on the environmental feedback during task executions. Experimental results on the NICOL robot for two classes of long-horizon everyday tasks showcase the superior performance of the LABOR agent compared to the Baseline agent without guiding prompts.
Limitations and Future Work. Similar to other LLM-based methods, the LABOR agent is dependent on the performance of the LLM like the GPT-4o model, requiring its reasoning and understanding of the correlations of primitive skills and task solutions. Besides, since the agent is currently solely running in the simulated environment instead of the real world, some potential differences in environmental dynamics might not be sufficiently analyzed or considered, like self-collision detection between the hands, which brings some unknowns in transferring the agent. In future work, the LLM can be further used to guide the learning of versatile skills in controlling humanoid hands [31, 35], which can endow the LABOR agent with fine-grained manipulation skills for everyday objects with various shapes. Moreover, vision-based foundation models [36, 37] can enhance the scene understanding capabilities of the LABOR agent when dealing with real-world tasks. As skills expand, the LABOR agent should be able to control bimanual robots to perform more complex long-horizon tasks.
@@ VI Acknowledgment
The authors would like to thank OpenAI and their Researcher Access Program for generously providing GPT-4o API tokens support, and Jan-Gerrit Habekost and Lennart Clasmeier for their work and support for the NICOL robot in the simulation.